# Un-Straightening Generative AI: How Queer Artists Surface and Challenge the Normativity of Generative AI Models

### v1

@@ Abstract.
Queer people are often discussed as targets of bias, harm, or discrimination in research on generative AI. However, the specific ways that queer people engage with generative AI, and thus possible uses that support queer people, have yet to be explored. We conducted a workshop study with 13 queer artists, during which we gave participants access to GPT-4 and DALL-E 3 and facilitated group sensemaking activities. We found our participants struggled to use these models due to various normative values embedded in their designs, such as hyper-positivity and anti-sexuality. We describe various strategies our participants developed to overcome these models’ limitations and how, nevertheless, our participants found value in these highly-normative technologies. Drawing on queer feminist theory, we discuss implications for the conceptualization of ”state-of-the-art” models and consider how FAccT researchers might support queer alternatives.
@@
copyright: rightsretainedjournalyear: 2025ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
In April 2024, Wired magazine published an investigation into visual generative AI (GenAI) models titled “Here’s How Generative AI Depicts Queer People” (Rogers, 2024). Echoing prior scholarship (Gillespie, 2024; Bianchi et al., 2023), the article revealed that models like DALL-E 3, Midjourney, and Sora produce highly stereotypical representations of queer people (e.g., having purple hair). Despite these biases, queer artists were already finding ways to repurpose GenAI for political resistance. Stephen and Craig, the married duo behind The Rupublicans Project, used GenAI to create satirical images of anti-LGBTQ+ politicians in drag, raising funds for queer causes (Wiggins, 2023). However, such creative uses are frequently constrained by GenAI platforms’ avowedly apolitical usage policies (Stapleton et al., 2023). While some queer people have managed to leverage these technologies for activism, they often face the challenge of working with systems that were not designed for them.
@@
Prior work has largely focused on how GenAI models represent queer people, leaving significant gaps in understanding queer people’s lived experiences with these technologies. In particular, the experiences of queer artists — a group for whom art-making is deeply entwined with cultural and political identity — remain underexplored. Queer communities cultivate different aesthetic relationships to art than those in dominant cultures. Queer camp sensibility appreciates the artifice of failed seriousness (Sontag, 2018), such as a film so earnestly bad that it becomes good. Partially due to a lack of representation, queer people often read queer narratives into ostensibly straight media (Russo, 1987; Halberstam, 2011; Sedgwick, 1993); hence, the affinity for Judy Garland and Disney villains (Sharmin and Sattar, 2018; Maddison, 2000). At the same time, queer culture has had an immense impact upon the arts. The musical genres of disco, house, and hyperpop were formed in the crucible of queer nightlife (Glick, 2024; MacDonald, 2024). Prior research on the relationship between art and GenAI has typically focused on identifying ethical concerns (Goetze, 2024; Kawakami and Venkatagiri, 2024; Jiang et al., 2023) or building tools (Wang et al., 2024; Choi et al., 2024). However, the unique relationship between queer people, art, and GenAI remains under-explored.
@@
In this work, we conducted a mixed-method workshop study with 13 queer artists over the course of 5 weeks. During this time period, we gave participants unlimited access to GPT-4 and DALL-E 3 — logging their interaction data — and facilitated weekly group sensemaking activities. Our participants encountered numerous normative values in the design of these models, such as biases against sex. To overcome these limitations, our participants developed workarounds, such as using lexical variation to evade moderation and model chaining to refine prompts.
@@ 2. Related Work
In this section, we first draw connections between queer theory and critical studies of technology. We then ground our study in prior work on the often fraught relationships between AI, art, and queerness.
@@ 2.1. Straightening & Queering Technology
We draw on queer theory to conceptualize the normative values embedded in the design of technology as well as how individuals respond to these normativities. Queer theory is often used to study the norms associated with gender and sexuality (Butler, 1988). However, much like how feminist studies is not only applicable to the study of women’s experiences (D’ignazio and Klein, 2020), queer theory is not only applicable to the study of queer people. Rather, queer theory draws on the lived experiences of queer people to understand the world. For example, research on queer temporalities has challenged the normative life schedules associated with the nuclear family — such as monogamous marriage and biological procreation (Halberstam, 2003) — which impact anyone who falls outside these social expectations. As queer lives often transgress social norms, queer theory provides a lens through which to interrogate these dominant norms (Halberstam, 2011).
@@
Technologies often function as straightening devices, orienting users toward some actions and away from others (Suchman, 1987). Madeleine Akrich refers to these assumptions about technology uses/users as scripts (Akrich, 1992). Sometimes designers create scripts to work against uses/users, such as deliberately making consumer electronics hard to repair to encourage people to buy new devices (Mattern, 2018). Even when designers try to work with users, human-computer interaction can still break down when users fail or are unable to follow designers’ plans (Suchman, 1987). It is when these breakdowns occur that the values embedded in technology can become most apparent (Bowker and Star, 1999), such as one’s gender not fitting provided options on a webform (Spiel, 2021). The scripts that designers—intentionally or not—embed in technology are inherently political and tend to reify established social orders (Suchman, 1993) by prescribing certain normative uses/users and marginalizing others (Baumer and Brubaker, 2017).
@@
Despite designers’ efforts to control user behavior, individuals need not abide by designers’ scripts (Akrich, 1992; Mackay and Gillespie, 1992). Users have agency over if or how they choose to take up technology, a process known as appropriation (Mackay and Gillespie, 1992). Sometimes, this takes the form of users working against designers’ intentions. For example, there are online forums dedicated to helping people fix devices that designers did not want to be repairable (Mattern, 2024). Likewise, people use lexical variation to evade algorithmic content moderation when posting on social media (Chancellor et al., 2016; Steen et al., 2023). Individuals may also leverage technologies in ways that designers never imagined, such as children using delicate Wi-Fi antennae as laptop handles (Ames, 2019). Drawing on Ahmed, these moments when users go off-script can be characterized as forms of ”queer use” (Ahmed, 2019). In this work, we contend with how our participants negotiate between the straightness of GenAI models and their own queer uses.
@@
There are numerous ways that individuals and collectives use algorithmic systems queerly. The grassroots audit of racial biases in Twitter’s image cropping algorithm required using the algorithm in an unintended way: to understand the algorithm itself. This queer use revealed straightening devices that, quite literally, oriented users toward whiteness (Shen et al., 2021). Evading algorithmic content moderation—be it in the context of social media or GenAI—is also a form of queer use because doing so requires deliberately subverting developers’ intentions (Chancellor et al., 2016; Steen et al., 2023). That being said, queer uses of algorithmic systems encompass more than resisting the technologies themselves, such as our earlier mention of queer artists using GenAI to resist anti-LGBTQ+ politicians (Stapleton et al., 2023). As mentioned above, Ahmed defines queer uses as both using things in ways they were not intended or when things are used by those who are not intended (Ahmed, 2019). Next section, we summarize prior work on two groups often under-considered in the design of AI systems: queer people and artists.
@@ 2.2. Queers, Artists, AI
Queer communities experience discrimination from algorithmic systems. Hate speech detection algorithms have been shown to be biased against the ways that some queer people speak (Thiago et al., 2021), similar to the biases against African American English in hate speech detection (Sap et al., 2019). Automated gender recognition algorithms harm transgender people by both failing to account for non-binary gender identities and reinforcing gender essentialist ideologies (Scheuerman et al., 2019). Queer people have to navigate online targeted advertising (Sampson et al., 2023) and social media algorithms (Simpson and Semaan, 2021) that amplify the most heteronormative representations of queerness, especially harming queer people of color. Moreover, the algorithmic moderation of nudity on social media has substantially harmed transgender communities (Mayworm et al., 2024). At the same time, queer people have resisted harmful algorithmic systems by, for instance, conducting collective audits to bring attention to algorithmic discrimination (Shen et al., 2021). Within the AI research community itself, the organization Queer in AI has worked to highlight issues like these, support queer researchers, and advocate for policy changes (Queerinai et al., 2023).
@@
The relationship between artists and GenAI is highly fraught. One major concern is the lack of consent from artists whose work was scraped to train GenAI models (Reisner and Gilbertson, 2024; Reisner, 2023a, b). Artists have fought back against unauthorized scraping by organizing grassroots data poisoning campaigns (Vincent et al., 2021), such as the online fan fiction community organizing a sexually explicit write-a-thon (Silberling, 2023). Computing researchers have also developed imperceptible pre-processing tools to help artists protect images they share online. In particular, Glaze helps artists prevent their style from being replicated by GenAI models (Shan et al., 2023) and Nightshade poisons text-to-image training data (Shan et al., 2024). Companies are also increasingly trying to use GenAI to replace artists’ jobs (Jiang et al., 2023; Kawakami and Venkatagiri, 2024; Merchant, 2024), a major concern in the 2023 Hollywood Writers Strike (Coyle, 2023).
At the same time, researchers have explored how artists use AI. Some artists use AI in their practices as tools to critique technology or raise matters of concern (Hemment et al., 2023; Caramiaux and Fdili Alaoui, 2022; Walker et al., 2023). Others have found that artists enjoy the glitches (Chang et al., 2023), uncertainty (Sivertsen et al., 2024), and surprises (Caramiaux and Fdili Alaoui, 2022) of working with stochastic AI models. That being said, there is a difference between research on those who may use AI in the process of their art making and those who identify as AI artists (Caramiaux and Fdili Alaoui, 2022). Recent developments in GenAI technology have led to an increased focus on the latter (Chang et al., 2023; Sanchez, 2023).
Those using GenAI in their artistic practices often have to contend with the biases embedded in these models (Jiang and Ha, 2020; Mim et al., 2024; Mirowski et al., 2024). Mirowski et al. found that some comedians struggled to write material using LLMs because content moderation systems inadvertently suppress jokes by members of marginalized groups about their own experiences (Mirowski et al., 2024). This work parallels hate speech scholarship that emphasizes the need for considering speaker identities when classifying the appropriateness of an utterance (Zhou et al., 2023). We build on Mirowski et al.’s work regarding marginalized communities generally by specifically focusing on queer artists’ experiences interacting with GenAI.
@@ 3. Methods
To understand queer artists’ experiences using GenAI, we gave a cohort of 13 artists access to GenAI models DALL-E 3 and GPT-4 and conducted a total of 10 remote workshops over Zoom during a 5-week period. Our workshops facilitated group discussion and deliberation surrounding Generative AI models, paralleling prior work that has used workshops to build critical consciousness around data (Markham, 2021) and support public deliberation over civic technology (Boehner and DiSalvo, 2016). Every week, we held two one-hour workshops with identical prompts and scaffolding to accommodate for timezone differences and schedule variability. Our participants could choose to attend either the morning or the evening session, but not both. In our findings, we use ‘AM’ or ‘PM’ to distinguish between workshops in the same week, such as ‘W5 AM.’
@@
Theme Pre-Work Instructions Workshop Description
W1 Onboarding Slide: Introduce your art and feelings about GenAI Intros, Onboarding
W2 Attitudes Toward GenAI
@@
Slide: Write a letter to an artist whose work was used to build the GenAI models in this study Share Back, Discussion
W3 Experiences Using
Current Models
Slide: Represent your experiences using or exploring GenAI during this study Share Back, Discussion
W4 Imagining Alternative
Futures
Slide: Represent how you would want to make, not make, or break GenAI Share Back, Discussion
W5 Synthesis No Slide: Individually reflect Discussion, Offboarding
@@
We gave our participants access to GPT-4 and DALL-E during our study through a website we built resembling common chat and image generation interfaces. This website acted as a “technology probe,” a technology deployed to collect use data in situ as well as inspire users to reflect on their technological needs (Hutchinson et al., 2003). For GPT-4, we logged each conversation’s system prompt, participant prompts, and model responses. For DALL-E 3, we logged every prompt and whether the model responded with an image or a content moderation error. Due to space limitations, we did not store the images generated by DALL-E 3. This observational data allowed us to understand how our participants were engaging with the provided models, helping situate issues raised by our participants in the workshops.
@@
Our participants engaged in a variety of artistic practices: creative writing (P2, P3), poetry (P4, P9, P10, P15), digital art (P3, P7, P8, P10), painting (P2, P4, P13), performance art (P14), textile art (P7, P11), and sculpture (P1, P5). While we did not ask participants directly about their economic relationships to art, throughout the study we learned that some teach art (P5, P14) and have created commissioned works (P1, P11). Others engaged in art making that is not intended to be commodified. For example, the fan fiction community (P2, P3) has strong norms against selling one’s work (Fiesler and Bruckman, 2019). Some participants also worked in artistic fields, such as design (P13) and architecture (P4). Our participants engaged in 142 unique GPT-4 chat conversations, in which they sent a total of 778 messages. Our participants attempted 2,092 DALL-E 3 prompts. The amount that our individual participants used GPT-4 and DALL-E 3 followed a power law distribution, with some using the models substantially more than others (Table 2).
@@
Table 2. Descriptive statistics of the number of GenAI prompts per participant.

### v2

@@ Abstract.
Queer people are often discussed as targets of bias, harm, or discrimination in generative AI research. However, the specific ways that queer people engage with generative AI, and thus possible uses that support queer people, have yet to be explored. We conducted a workshop study with 13 queer artists, during which we gave participants access to GPT-4 and DALL-E 3 and facilitated group sensemaking activities. Our participants struggled to use these models due to various normative values embedded in their designs, such as hyper-positivity and anti-sexuality. We describe various strategies our participants developed to overcome these models’ limitations and how, nevertheless, some found value in these highly-normative technologies. Drawing on queer feminist theory, we discuss implications for the conceptualization of ”state-of-the-art” models and consider how FAccT researchers might support queer alternatives.
@@
copyright: acmcopyrightjournalyear: 2025ccs: Human-centered computing Empirical studies in HCI
Figure 1. A slide created by participant P13 of 24 attempts to generate DALL-E 3 images of “intimacy.” The images are overlaid on a photo of two men kissing at a protest, leaving gaps for the five times their prompt was denied by DALL-E 3’s moderation system. This slide critiques the heteronormativity of DALL-E 3, imagining queer intimacy may lie in what cannot be generated.
[…]
A workshop slide created by one of our participants representing their 24 repeated attempts to generate images of “intimacy” using DALL-E 3. On the left of the slide is the participants’ ID, text reading ”24 attempts at intimacy,” and a quote attributed to Barbara Kruger reading ”You construct intricate rituals which allow you to touch the skin of other men.” One the right of this text is a rectangular 4 by 6 image collage. 19 AI-generated images showing heteronormative couples and people holding hands are overlaid on a photo of two men kissing, leaving gaps with error messages through which the men kissing behind the collage can be seen.
@@ 1. Introduction
In April 2024, Wired magazine published an investigation into visual generative AI (GenAI) models titled “Here’s How Generative AI Depicts Queer People” (Rogers, 2024). Echoing prior scholarship (Gillespie, 2024; Bianchi et al., 2023), the article revealed that models like DALL-E 3, Midjourney, and Sora produce highly stereotypical representations of queer people (e.g., having purple hair). Despite these biases, queer artists were already finding ways to repurpose GenAI for political resistance. Stephen and Craig, the married duo behind The Rupublicans Project, used GenAI to create satirical images of anti-LGBTQ+ politicians in drag, raising funds for queer causes (Wiggins, 2023). However, such uses are frequently constrained by GenAI platforms’ avowedly apolitical usage policies (Stapleton et al., 2023). While some queer people have managed to leverage these technologies for activism, they often face the challenge of working with systems that were not designed for them.
@@
Prior work has largely focused on how GenAI models represent queer people, leaving significant gaps in understanding queer people’s lived experiences with these technologies. In particular, the experiences of queer artists — a group for whom art-making is deeply entwined with cultural and political identity — remain underexplored. Queer communities cultivate different aesthetic relationships to art than those in dominant cultures. Queer camp sensibility appreciates the artifice of failed seriousness (Sontag, 2018), such as a film so earnestly bad that it becomes good. Partially due to a lack of representation, queer people often read queer narratives into ostensibly straight media (Russo, 1987; Halberstam, 2011; Sedgwick, 1993); hence, the affinity for Judy Garland and Disney villains (Sharmin and Sattar, 2018; Maddison, 2000). Queer culture has also had an immense impact upon the arts. The musical genres of disco, house, and hyperpop were formed in the crucible of queer nightlife (Glick, 2024; MacDonald, 2024). Queer art can also refer to non-normative aesthetic styles (Lorenz, 2012). Art made by queer people or in queer styles is often denigrated by social conservatives. Nazi Germany propagandists referred to more-abstract Modern Art as “Degenerate Art,” associating the style with Jewishness, communism and homosexuality (Smee, 2025). In the USA, queer art is often the target of government censorship, ranging from attempts to ban Allen Ginsberg’s poetry in the 1950s (Morgan, 2021) to the proliferation drag bans at the time of our writing (Marwick et al., 2024).
@@
Prior research on art and GenAI has typically focused on identifying ethical concerns (Goetze, 2024; Kawakami and Venkatagiri, 2024; Jiang et al., 2023) or building tools (Wang et al., 2024; Choi et al., 2024). However, the unique relationship between queer people, art, and GenAI remains under-explored. In this work, we conducted a mixed-method workshop study with 13 queer artists over the course of 5 weeks. During this time period, we gave participants unlimited access to GPT-4 and DALL-E 3 — logging their interaction data — and facilitated weekly group sensemaking activities. Our participants encountered numerous normative values in the design of these models, such as biases against sex. To overcome these limitations, our participants developed workarounds, such as using lexical variation to evade moderation and model chaining to refine prompts.
@@ 2. Related Work
In this section, we draw connections between queer theory and critical studies of technology. Then, we ground our study in prior work on the often fraught relationships between AI, art, and queerness.
@@ 2.1. Straightening & Queering Technology
We draw on queer theory to conceptualize the normative values embedded in the design of technology as well as how individuals respond to these normativities. Queer theory is often used to study the norms associated with gender and sexuality (Butler, 1988). Of course, not all queer people necessarily resist dominant social norms and queer theory is not intended to represent the experiences of all queer people. Instead, queer theory draws on the lived experiences of queer people to understand the world. Much like how feminism is not only applicable to the study of women’s experiences (D’ignazio and Klein, 2020), queer theory is also not only applicable to the study of queer people. For example, research on queer temporalities has challenged the normative life schedules associated with the nuclear family — such as monogamous marriage and biological procreation (Halberstam, 2003) — which impact anyone who falls outside these social expectations. As queer lives often transgress social norms, queer theory provides a lens through which to interrogate these dominant norms (Halberstam, 2011). Similar to our use of queer theory, prior FAccT research has used queer theory to critique the normativity (e.g., thinness and whiteness) of queer people’s experiences with targeted advertising (Sampson et al., 2023).
@@
Technologies often function as straightening devices, orienting users toward some actions and away from others (Suchman, 1987). Madeleine Akrich refers to these assumptions about technology uses/users as scripts (Akrich, 1992). Sometimes designers create scripts to work against uses/users, such as making electronics hard to repair to encourage people to buy new devices (Mattern, 2018). Even when designers try to work with users, human-computer interaction can still break down when users fail or are unable to follow designers’ plans (Suchman, 1987). The values embedded in technology can become most apparent when these breakdowns occur (Bowker and Star, 1999), such as one’s gender not fitting the options on a webform (Spiel, 2021). The scripts that designers—intentionally or not—embed in technology are inherently political and tend to reify established social orders (Suchman, 1993) by prescribing certain normative uses/users and marginalizing others (Baumer and Brubaker, 2017).
@@
Despite designers’ efforts to control user behavior, individuals need not abide by designers’ scripts (Akrich, 1992; Mackay and Gillespie, 1992). Users have agency over if or how they choose to take up technology, a process known as appropriation (Mackay and Gillespie, 1992). Sometimes, this takes the form of users working against designers’ intentions, such as the online forums dedicated to helping people fix devices that designers did not want to be repairable (Mattern, 2024). Likewise, people use lexical variation to evade algorithmic content moderation when posting on social media (Chancellor et al., 2016; Steen et al., 2023). Individuals may also leverage technologies in ways that designers never imagined, such as children using delicate Wi-Fi antennae as laptop handles (Ames, 2019). Drawing on Ahmed, these moments when users go off-script can be characterized as forms of ”queer use” (Ahmed, 2019). In this work, we contend with how our participants negotiate between the straightness of GenAI models and their own queer uses.
@@
There are numerous ways that individuals and collectives use algorithmic systems queerly. The grassroots audit of racial biases in Twitter’s image cropping algorithm required using the algorithm in an unintended way: to understand the algorithm itself. This queer use revealed straightening devices that, quite literally, oriented users toward whiteness (Shen et al., 2021). Evading algorithmic moderation — in the context of social media or GenAI — is also a form of queer use because doing so requires deliberately subverting developers’ intentions (Chancellor et al., 2016; Steen et al., 2023; Calhoun and Fawcett, 2023). That said, queer uses of algorithmic systems encompass more than resisting the technologies themselves, such as our earlier mention of queer artists using GenAI to resist anti-LGBTQ+ politicians (Stapleton et al., 2023). As mentioned above, Ahmed defines queer uses as both using things in ways they were not intended or when things are used by those who are not intended (Ahmed, 2019). Next section, we summarize prior work on two groups often under-considered in the design of AI systems: queer people and artists.
@@ 2.2. Queers, Artists, AI
Queer communities experience discrimination from algorithmic systems. Hate speech detection algorithms have been shown to be biased against the ways that some queer people speak (Thiago et al., 2021; Thylstrup and Talat, 2020), similar to the biases against African American English (Sap et al., 2019). Automated gender recognition algorithms harm transgender people by both failing to account for non-binary gender identities and reinforcing gender essentialist ideologies (Scheuerman et al., 2019). Queer people navigate online targeted advertising (Sampson et al., 2023) and social media algorithms (Simpson and Semaan, 2021) that amplify the most heteronormative representations of queerness, especially harming queer people of color. Moreover, the algorithmic moderation of nudity on social media has substantially harmed transgender communities (Mayworm et al., 2024). At the same time, queer people have resisted harmful algorithmic systems by, for instance, conducting collective audits to bring attention to algorithmic discrimination (Shen et al., 2021). Within the AI research community itself, the organization Queer in AI has worked to highlight issues like these, support queer researchers, and advocate for policy changes (Queerinai et al., 2023).
@@
The relationship between artists and GenAI is highly fraught. Companies are increasingly trying to use GenAI to replace artists’ jobs (Jiang et al., 2023; Kawakami and Venkatagiri, 2024; Merchant, 2024), a major concern in the 2023 Hollywood Writers Strike (Coyle, 2023). At the same time, artists’ work was scraped to train these models without their consent (Reisner and Gilbertson, 2024; Reisner, 2023a, b). Artists have fought against unauthorized scraping by organizing grassroots data poisoning campaigns (Vincent et al., 2021), such as the online fan fiction community organizing a sexually explicit write-a-thon (Silberling, 2023). Researchers have also developed imperceptible pre-processing tools to help artists protect images they share online. Glaze helps artists prevent their style from being replicated by GenAI models (Shan et al., 2023) and Nightshade poisons text-to-image training data (Shan et al., 2024).
Researchers have also explored how artists use AI. Some artists use AI in their practices as tools to critique technology or raise matters of concern (Hemment et al., 2023; Caramiaux and Fdili Alaoui, 2022; Walker et al., 2023). Others have found that artists enjoy the glitches (Chang et al., 2023), uncertainty (Sivertsen et al., 2024), and surprises (Caramiaux and Fdili Alaoui, 2022) of working with stochastic AI models. That being said, there is a difference between research on those who may use AI in the process of their art making and those who identify as AI artists (Caramiaux and Fdili Alaoui, 2022). Recent developments in GenAI technology have led to an increased focus on the latter (Chang et al., 2023; Sanchez, 2023). Bennett et al.’s research on how creatives with disability choose (not) to use AI demonstrates how artists negotiations potential benefits, such as using AI to access new mediums, alongside ethical concerns (Bennett et al., 2024). For example, those using GenAI in their artistic practices have to contend with the biases embedded in these models (Jiang and Ha, 2020; Mim et al., 2024; Mirowski et al., 2024). Mirowski et al. found that some comedians struggled to write material using LLMs because content moderation systems inadvertently suppress jokes by members of marginalized groups about their own experiences (Mirowski et al., 2024). This work parallels hate speech scholarship that emphasizes the need for considering speaker identities when classifying the appropriateness of an utterance (Zhou et al., 2023). We build on Mirowski et al.’s work regarding marginalized communities generally by specifically focusing on queer artists’ experiences interacting with GenAI.
@@ 3. Methods
To understand queer artists’ experiences using GenAI, we gave a cohort of 13 artists access to DALL-E 3 and GPT-4 and conducted a total of 10 remote workshops over Zoom during a 5-week period. Our workshops facilitated group discussion and deliberation surrounding GenAI models, paralleling prior work that has used workshops to build critical consciousness around data (Markham, 2021) and support public deliberation over civic technology (Boehner and DiSalvo, 2016). Every week, we held two one-hour workshops with identical prompts (Table 1) and scaffolding (i.e., onboarding information) to accommodate for timezone differences and schedule variability. Our participants could choose to attend either the morning or the evening session, but not both. In our findings, we use ‘AM’ or ‘PM’ to distinguish between workshops in the same week, such as ‘W5 AM.’
@@
Theme Pre-Work Instructions
W1 Onboarding Slide: Introduce your art and feelings about GenAI
W2 Attitudes
Toward GenAI
@@
Slide: Write a letter to an artist whose work was used to build the GenAI models in this study
W3 Experiences Using Models Slide: Represent your experiences using or exploring GenAI during this study
W4 Imagining
Alternatives
Slide: Represent how you would want to make, not make, or break GenAI
W5 Synthesis No Slide: Individually reflect
@@
We gave our participants access to GPT-4 and DALL-E during our study through a website we built resembling common chat and image generation interfaces. We deployed these models using Microsoft’s enterprise Azure OpenAI cloud service in February of 2024. For DALL-E 3 we used the default parameters for size (1024x1024), style (vivid), and quality (standard). For GPT-4 we used version 0613 with the following parameters: temperature (0.7), max_tokens (800), top_p (0.95), frequency_penalty (0), presence_penalty (0). Our study site acted as a “technology probe,” one deployed to collect use data in situ and inspire users to reflect on their technological needs (Hutchinson et al., 2003). For GPT-4, we logged each conversation’s system prompt, participant prompts, and model responses. For DALL-E 3, we logged every prompt and whether the model responded with an image or a content moderation error. Due to space limitations, we did not store the DALL-E 3 generated images. This observational data allowed us to understand how our participants were engaging with the models, helping situate issues raised in the workshops.
@@
Our participants engaged in a variety of artistic practices: creative writing (P2, P3), poetry (P4, P9, P10, P15), digital art (P3, P7, P8, P10), painting (P2, P4, P13), performance art (P14), textile art (P7, P11), and sculpture (P1, P5). While we did not ask participants directly about their economic relationships to art, throughout the study we learned that some teach art (P5, P14) and have created commissioned works (P1, P11). Others engaged in art making that is not intended to be commodified. For example, the fan fiction community (P2, P3) has strong norms against selling one’s work (Fiesler and Bruckman, 2019). Some participants also worked in artistic fields, such as design (P13) and architecture (P4). To protect participant privacy, we purposefully did not require participants to share standardized demographic information about themselves. We also sought to empower participants to choose how they wished to bring aspects of their identities into the study. Therefore, beyond our inclusion criteria that participants identify as queer artists, we only requested participants’ pronouns and optionally asked, “Are there any identities you hold that you wish to share?” In the work, we refer to each participant according to these pronouns and discuss aspects of participants identities within the the findings. In aggregate, at least five of our participants identified as transgender or non-binary. While not all participants shared their racial identities, one identified as Black, one as Native American, and one as Chinese. To the best of our knowledge, all of our participants were adults between the ages of 20 and 50 and all lived in urban areas of the Midwest or East Coast of the USA. Our participants engaged in 142 unique GPT-4 chat conversations, in which they sent a total of 778 messages. Our participants attempted 2,092 DALL-E 3 prompts. The amount that our individual participants used GPT-4 and DALL-E 3 followed a power law distribution, with some using the models substantially more than others (Table 2).
@@
Table 2. Description of participant GenAI use